Pixel:

What is the equation of a line that is parallel to 2x+3y=3 and passes through the point (3, −4) ?

RickandMorty:

google it

Pixel:

i did

RickandMorty:

the googled it wrong

RickandMorty:

google always has an answer you just have to google it right

AP:

Start by writing the equation in slope-intercept form. (y = mx + b)

Pixel:

3y=2x+3

AP:

Actually, you would need to solve the equation like you're solving for 'y' like this 

Pixel:

oof

Pixel:

so the slope is -2/3

AP:

Yes, and parallel lines share the same slope, correct

Pixel:

Okay, but what is the Y itercept

Pixel:

input the coordinates?

AP:

Yes, plug the coordinates into the equation and solve for 'b'.

Pixel:

Im confused

AP:

The (x, y) coordinates: (3, -4). Substitute them into your equation y=-2/3x + b 

Pixel:

so y =-2/3 - 2

AP:

Now, you know your slope and your y-intercept

AP:

Correct.

AP:

You can always check your work like this: https://www.desmos.com/calculator/avncdadq0e

Pixel:

i tagged you in another question i need help onn